Use the following data for questions 25 thru 31:

Cox Engineering performs cement core tests in its laboratory. The following standards have been set for each core test

performed based on planned activity of 1,800 core tests:

Standard Hours or

Quantity

Standard Price or

Rate

Price Per

Unit

Direct materials ........................................

3 pounds

$0.75 per pound

$2.25

Direct labor...............................................

0.4 hours

$12 per hour

$4.80

Variable manufacturing overhead ............

0.4 hours

$9 per hour

$3.60

Fixed manufacturing overhead.............

$6,800

During March, the laboratory performed 2,000 core tests. On March 1 no direct materials (sand) were on hand. Variable

manufacturing overhead is assigned to core tests on the basis of standard direct labor-hours. The following events

occurred during March:

• 7,200 pounds of sand were purchased at a cost of $6,120.

• 7,200 pounds of sand were used for core tests.

• 840 actual direct labor-hours were worked at a cost of $8,610.

• Actual variable manufacturing overhead incurred was $7,240.

• Actual fixed manufacturing overhead incurred was $6,500.

25.

The materials price variance for March is:

A. $720 unfavorable

B. $720 favorable

C. $281 unfavorable

D. $281 favorable

26.

The materials quantity variance for March is:

A. $900 favorable

B. $1,950 favorable

C. $1,950 unfavorable D. $900 unfavorable

27.

The labor rate variance for March is:

A. $4,578 unfavorable B. $1,470 unfavorable C. $4,578 favorable D. $1,470 favorable

28.

The labor efficiency variance for March is:

A. $480 favorable

B. $480 unfavorable C. $192 favorable

D. $192 unfavorable

29.

The variable manufacturing overhead price variance for March is:

A. $320 unfavorable

B. $320 favorable

C. $40 unfavorable

D. $40 favorable

30.

The variable manufacturing overhead efficiency variance for March is:

A. $360 unfavorable

B. $360 favorable

C. $40 unfavorable

D. $40 favorable

31.

The total fixed manufacturing overhead variance for March is:

A. $300 unfavorable

B. $300 favorable

C. $40 unfavorable

D. $40 favorable
Answer #1
Material Labour Variable O/H Fixed O/H
Standard Standard Standard Standard
Qty Price Total p.u Hours Rate Total p.u Hours Rate Total p.u
3 0.75 2.25 0.4 12 4.8 0.4 9 3.6 6800
Standard (1800 Tests) Standard (1800 Tests) Standard (1800 Tests) Standard (1800 Tests)
Qty Price Total p.u Hours Rate Total p.u Hours Rate Total p.u
5400 0.75 4050 720 12 8640 720 9 6480 6800
Actual (2000 Tests) Actual (2000 Tests) Actual (2000 Tests) Actual (2000 Tests)
Qty Price Total p.u Hours Rate Total p.u Hours Rate Total p.u
7200 0.85 6120 840 10.25 8610 840 8.62 7240 6500
Material Price Variance =(S.P-A.P)*AQ Labour Rate Variance =(S.R-A.R)*AH V O/H Price Variance =(S.R-A.R)*AH Total fixed manufacturing overhead variance
=(0.75-0.85)*7200 -720 =(12-10.25)*840 1470 =(9-8.62)*840 319.2 =Standard Fixed O/H-Actual Fixed O/H
=6800-6500
Material Qty Variance =(S.Q-A.Q)*SP Labour Efficiency Variance =(S.H-A.H)*SR V O/H Efficiency Variance =(S.H-A.H)*SR 300
=(3*2000-7200)*0.75 -900 =(0.4*2000-840)*12 -480 =(0.4*2000-840)*9 -360

25.  The materials price variance for March is: A. $720 unfavorable

26.  The materials quantity variance for March is:   D. $900 unfavorable

27.  The labor rate variance for March is: D. $1,470 favorable

28.  The labor efficiency variance for March is: B. $480 unfavorable

29.  The variable manufacturing overhead price variance for March is: B. $320 favorable

30.  The variable manufacturing overhead efficiency variance for March is: A. $360 unfavorable

31.  The total fixed manufacturing overhead variance for March is:  B. $300 favorable

(Kindly Rate Positively, if found helpful.)